There is a 3 kg mass being swung around by a rope in a circular motion. The rope holding the mass is 0.75 meters long and can withstand a max tension of 20N. a) What is maximum centripetal acceleration before this rope breaks b) What is the maximum speed of the mass at this point.
Answer #1

a)

a = F/m = 20 / 3 = 6.667 m/s^2

b)

a = v^2/r

6.667 = v^2/ 0.75

v = 2.237 m/s

====

Comment before rate in case any doubt, will reply for sure.. goodluck